You are standing at the edge of a cliff that drops straight down to a river below. You reach out over the cliff and toss a stone straight up in the air. The stpne passes you on its way down, 5.00 s after it was thrown. What is the speed of the rock just before it reaches the water 27.0 m below the point where the stone left your hand? Ignore air resistance. Express your answer with the appropriate units.
Answer #1

Initial velocity, u = t g/2 = 5 x 9.8/2 = 24.5 m/s

0.5 mv^2 - 0.5 mu^2 = mgh

V = sqrt[(2 x 9.8 x 27) + (24.5^2)]

V = 33.61 m/s

Comment in case any doubt please rate my answer ....